is y squared = x+5 a relation or a function

Answer by jim_thompson5910(35256)

y^2 = x+5

y = sqrt(x+5) or y = -sqrt(x+5)

so it is NOT a function (since most of the x values map to more than one y value, eg: x = -1 maps to both y = 2 and y = -2)